# TokenSHAP: Interpreting Large Language Models with Monte Carlo Shapley Value Estimation

### v1

@@ 3.3 Monte Carlo Shapley Estimation
We adapt the Monte Carlo sampling approach proposed by Castro et al. [16] to the context of text inputs. For each token $x_{i}$, we estimate its Shapley value $\phi_{i}$ as follows:
@@
    Randomly select a subset size $k\in\{0,...,n-1\}$
@@
    Randomly select a subset $S$ of size $k$ that does not include $i$
@@
    Compute the marginal contribution: $v(S\cup\{i\})-v(S)$
@@
    Repeat steps 1-3 $M$ times and average the results
This estimator is unbiased and converges to the true Shapley value as $M$ increases, ensuring the robustness and reliability of the importance scores generated. Importantly, in our adaptation, we methodically ensure that each token $x_{i}$ is excluded precisely once in the subsets considered for each calculation. This consistent exclusion of each token is crucial, as it provides a stable calculation base, ensuring that the estimations are not only unbiased but also systematically verifiable across different textual inputs. By doing so, we maintain methodological integrity, allowing for a robust basis of comparison and significance testing among tokens’ contributions to the model output.
@@
9:$N\leftarrow\lfloor(2^{n}-1)\cdot r\rfloor$triangleright Number of sampled combinations
10:$S\leftarrow$ Random sample of $N-n$ combinations excluding $E$
11:$C\leftarrow E\cup S$triangleright All combinations to process
12:for each combination $c$ in $C$ do
13:     Get model response $R_{c}$ for $c$
14:     Calculate cosine similarity $\text{sim}(b,R_{c})$
15:end for
16:for $i=1$ to $n$ do
17:     $\text{with}_{i}\leftarrow$ average similarity of combinations including $x_{i}$
18:     $\text{without}_{i}\leftarrow$ average similarity of combinations excluding $x_{i}$
19:     $\phi_{i}\leftarrow\text{with}_{i}-\text{without}_{i}$
20:end for
21:Normalize ${\phi_{1},\ldots,\phi_{n}}$
22:return ${\phi_{1},\ldots,\phi_{n}}$
Algorithm 1 provides a high-level overview of the TokenSHAP method. It iterates over each sample, performing Monte Carlo sampling to estimate its Shapley value. The algorithm efficiently computes marginal contributions and averages them to produce the final Shapley values for all tokens in the input prompt.
3.4 Value Function
3.5 Model Interaction
3.6 Shapley Value Computation
The estimated Shapley value for token $x_{i}$ is computed as:
$\phi_{i}=\frac{1}{M}\sum_{m=1}^{M}[v(S_{m}\cup\{i\})-v(S_{m})]$ (2)
where $S_{m}$ are the randomly sampled subsets, and $M$ is the number of Monte Carlo iterations. This aggregation of marginal contributions over multiple samples ensures a robust and reliable estimation of token importance.
@@ 4.2.3 Results and Analysis
Figure 6 presents the results, indicating significant differences in approximation accuracy depending on the presence of the first-order omission condition. With this condition, even at lower sampling ratios such as 0.1 and 0.0, the similarity between the estimated and true Shapley values remains relatively stable or shows a slight increase. This suggests that smaller sampling ratios, without including the first-order omission, primarily introduce noise, thus negatively impacting performance. Conversely, starting from a sampling ratio of 0.4 and higher, the noise diminishes, and the accuracy of the approximations exceeds those observed at lower ratios.

### v2

@@ 3.3 Monte Carlo Shapley Estimation
We adapt the Monte Carlo sampling approach to the context of text inputs. For each token $x_{i}$, we estimate its Shapley value $\phi_{i}$ as follows:
@@
    Generate a set of combinations that includes:
    1. (a)
      All combinations where $x_{i}$ is the only token removed (essential combinations)
    2. (b)
      A random sample of other combinations based on a specified sampling ratio
@@
    For each combination:
    1. (a)
      Generate the model’s response
    2. (b)
      Calculate the cosine similarity between this response and the full prompt response
@@
    Compute the average similarity for combinations with and without $x_{i}$
@@
    Calculate $\phi_{i}$ as the difference between these averages
This Monte Carlo estimation approach ensures a balance between computational efficiency and estimation accuracy. The use of essential combinations alongside random samples provides a robust basis for estimating Shapley values, even with a relatively small number of samples.
3.4 Value Function
3.5 Model Interaction
3.6 Shapley Value Computation
The estimated Shapley value for token $x_{i}$ is computed as:
[…]
This difference in average similarities provides a measure of the token’s importance to the model’s output. The final Shapley values are normalized to ensure comparability across different inputs and models.
@@ 3.6 Shapley Value Computation
9:$N\leftarrow\min(n,\lfloor(2^{n}-1)\cdot r\rfloor)$triangleright Number of sampled combinations
10:if $N<n$ then
11:     $C\leftarrow E$triangleright Use only first-order samples
12:else
13:     $S\leftarrow$ Random sample of $N-n$ combinations excluding $E$
14:     $C\leftarrow E\cup S$triangleright All combinations to process
15:end if
16:for each combination $c$ in $C$ do
17:     Get model response $R_{c}$ for $c$
18:     Calculate cosine similarity $\text{sim}(b,R_{c})$
19:end for
20:for $i=1$ to $n$ do
21:     $\text{with}_{i}\leftarrow$ average similarity of combinations including $x_{i}$
22:     $\text{without}_{i}\leftarrow$ average similarity of combinations excluding $x_{i}$
23:     $\phi_{i}\leftarrow\text{with}_{i}-\text{without}_{i}$
24:end for
25:Normalize ${\phi_{1},\ldots,\phi_{n}}$
26:return ${\phi_{1},\ldots,\phi_{n}}$
@@ 4.2.3 Results and Analysis
Figure 6 presents the results, demonstrating significant differences in approximation accuracy depending on the presence of the first-order omission condition. The sampling ratio plays a crucial role in determining the number of combinations considered beyond the essential first-order samples. It’s important to note:
  • A sampling ratio of 0.0 means only the essential first-order combinations are used, with no additional combinations.
  • At low sampling ratios (e.g., 0.1), the number of additional combinations is often negligible, especially for smaller input sizes, effectively yielding results similar to the 0.0 ratio.
  • As the sampling ratio increases (e.g., 0.2), a small number of additional combinations are included. However, this limited increase may introduce some noise without significantly improving accuracy.
  • With the first-order omission condition present, even at lower sampling ratios (0.1 and 0.0), the similarity between estimated and true Shapley values remains relatively stable or shows a slight increase.
  • Starting from a sampling ratio of 0.4 and higher, the increased number of combinations begins to reduce noise, and the accuracy of the approximations exceeds those observed at lower ratios.